an achitect designs a rectangular flower garden such that the width is exactly two-thirds of the length. if 250 feet of antique picket fencing are to be used to enclose the garden,find the dimensions of the garden
Answer by checkley77(12844)

W=2L/3
2L+2W=250
2L+2(2L/3)=250
2L+4L/3=250
(2L*3+4L)/3=250
(10L+2L)/3=250
10L=3*250
10L=750
L=750/10
L=75 ANS. FOR THE LENGTH.
W=2*75/3
W=150/3
W=50 ANS. FOR THE WIDTH.
PROOF:
2*75+2*50=250
150+100=250
250=250